OpenStudy (anonymous):

If 3 suspects who committed a burglary and 9 innocent persons are lined up what is the probability that a witness who is not sure and has to pick 3 persons will a. pick the three suspects by chance? b. pick three innocent people by chance? If 3 suspects who committed a burglary and 9 innocent persons are lined up what is the probability that a witness who is not sure and has to pick 3 persons will a. pick the three suspects by chance? b. pick three innocent people by chance? @Mathematics

OpenStudy (zarkon):

a)\[\frac{1}{_{12}C_{3}}\] b)\[\frac{_9C_3}{_{12}C_3}\]

OpenStudy (anonymous):

okay, i kinda don't get the C here. Could you perhaps explain how you arrived at the answer and what is it simplified?

OpenStudy (zarkon):

\[_nC_r={n\choose r}=\frac{n!}{r!(n-r)!}\]